OpenStudy (anonymous):

What is 8.75% written as a fraction in lowest terms?

OpenStudy (anonymous):

@pooja195

OpenStudy (just_one_last_goodbye):

Do you need help my friend? :)

OpenStudy (richie63):

7/80

OpenStudy (anonymous):

LOL yes plz!

OpenStudy (anonymous):

That's not right richie

OpenStudy (richie63):

medal pls

OpenStudy (just_one_last_goodbye):

Write the decimal number as a fraction (over 1) 8.75 = 8.75 / 1 Multiplying by 1 to eliminate 2 decimal places we multiply top and bottom by 2 10's Numerator (N) N = 8.75 × 10 × 10 = 875 Denominator (D) D = 1 × 10 × 10 = 100 N / D = 875 / 100 Simplifying our fraction = 875/100 = 35/4 = which makes the answer?

OpenStudy (richie63):

7/80 Make whole numbers by multiplying numerator and denominator by 100: 8.75% = 8.75/100 = 875/10000 Then reduce: 875/10000 reduced by common divisor 1/125 = 7/80ths.

OpenStudy (anonymous):

It makes the answer 7/80

OpenStudy (just_one_last_goodbye):

Let us remember @sydneymmelin the 8. means it's a whole :) so the transfer from the decimal to the fraction, the fraction should have 8 as a whole and not part of the fraction